Pearl Products Limited of Shenzhen, China, manufactures and distributes toys throughout South East Asia. Three cubic centimeters (cc) of solvent H300 are required to manufacture each unit of Supermix, one of the company’s products. The company now is planning raw materials needs for the third quarter, the quarter in which peak sales of Supermix occur. To keep production and sales moving smoothly, the company has the following inventory requirements:

  1. The finished goods inventory on hand at the end of each month must equal 2,000 units of Supermix plus 25% of the next month’s sales. The finished goods inventory on June 30 is budgeted to be 11,250 units.

  2. The raw materials inventory on hand at the end of each month must equal one-half of the following month’s production needs for raw materials. The raw materials inventory on June 30 is budgeted to be 57,375 cc of solvent H300.

  3. The company maintains no work in process inventories.

A monthly sales budget for Supermix for the third and fourth quarters of the year follows.

Budgeted Unit Sales
July 37,000
August 42,000
September 52,000
October 32,000
November 22,000
December 12,000

Required:

1. Prepare a production budget for Supermix for the months July, August, September, and October.

3. Prepare a direct materials budget showing the quantity of solvent H300 to be purchased for July, August, and September, and for the quarter in total.
Answer #1

Solution :

(1) Production Budget :

Particulars

July

August

September

October

(a) Budgeted Unit Sales

37,000 42,000 52,000 32,000

(b) Desired Inventory at the end

12,500 15,000 10,000 7,500

(c) Total Needs (a + b)

49,500 57,000 62,000 39,500

(d) Opening Inventory

11,250 12,500 15,000 10,000

(e) Required Production in Units (c - d)

38,250 44,500 47,000 29,500

Desired Inventory at the end of the month :

July : 2,000 + ( 42,000 * 0.25) = 12,500

Aug : 2,000 + ( 52,000 * 0.25) = 15,000

Sep : 2,000 + ( 32,000 * 0.25) = 10,000

Oct : 2,000 + ( 22,000 * 0.25) = 7,500

(2) Direct Material Budget :

Particulars

July

August

September

Third Quarter

(a) Required Production in Units

38,250 44,500 47,000 129,750

(b) Unit of Raw Material Required per Unit of Finished Goods

3 CC

3 CC

3 CC

3 CC

(c) Units of Raw Material Needed to Meet Production ( a * b)

114,750 133,500 141,000 389,250

(d) Desired Raw Material Units at the end of the month

66,750 70,500 44,250 44,250

(e) Total Units of Raw Material Needed (c +d )

181,500 204,000 185,250 433,500

(f) Opening Units of Raw Material

57,375 66,750 70,500 57,375

(g) Units of Raw Material to be Purchased ( e-f)

124,125 137,250 114,750

376,125

Desired Raw Material Units at the end of the month :

July : 133,500 * 0.50 = 66,750

Aug : 141,000 * 0.50 = 70,500

Sep : 29,500 * 3 * 0.50 = 44,250